OpenStudy (anonymous):

f(x) = x 2 + 1, find f(a + 1)

OpenStudy (firekat97):

since f(x) = x^2 + 1, to find f(a + 1), you want to substitute (a + 1) instead of x, on the right hand side and simplify that, so try that

OpenStudy (firekat97):

so you should get f(a + 1) = (a + 1)^2 + 1 now you should to simplify this